OpenStudy (anonymous):

Please help :) Describe the roots of the equation shown below. 64x^2-16x+1=0 A. There are two real, irrational roots. B. There is one real, double root. C. There are two real, rational roots. D. There are two complex roots.

OpenStudy (anonymous):

Do you have a y idea what it might be?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

WHy?

OpenStudy (anonymous):

have you solved?

OpenStudy (anonymous):

You can use quadratic equation to solve

OpenStudy (anonymous):

Yeah I know how to solve it but just not sure which one haha

OpenStudy (mertsj):

Factor it.

OpenStudy (anonymous):

both real and rational unless imaginary numbers were put into equation (Im assuming not so considering it wasnt mentioned in the question itself), and also one has negative coefficient, theres no need to even consider it. No need to solve

OpenStudy (mertsj):

\[64x^2-16x+1=(8x-1)(8x-1)=0\]

OpenStudy (mertsj):

\[8x-1=0 or 8x-1=0\]

OpenStudy (anonymous):

so 1/8

OpenStudy (mertsj):

\[x=\frac{1}{8}orx =\frac{1}{8}\]

OpenStudy (mertsj):

1 Double Root

OpenStudy (anonymous):

I'm so confused so it's not two real rational like what Reealcheyenne said?

OpenStudy (mertsj):

no

OpenStudy (mertsj):

There is one root which occurs twice. That is what we call a double root. It corresponds to a point of tangency of the graph to the x axis at that number.

OpenStudy (anonymous):

ohhh! alright thanks so much :)

OpenStudy (mertsj):

yw